OpenStudy (anonymous):

solve the seperable equation: x(dv/dx)= (1-4v^2)/3v im not sure how to get to the final answer

OpenStudy (across):

I already worked this one out for you...

OpenStudy (anonymous):

x(dv/dx)= (1-4v^2)/3v [3v/(1-4v^2)] dv =dx/x let u=1-4v^2, du=-8vdv -(3/8)du/u=dx/x -3/8 lnu =lnx +c -3/8 ln(1-4v^2) =lnx +c